OpenStudy (anonymous):

50x^2-15x-5x=0. I think i did this right, but I'm not sure. What are the solutions and how do you get them.

OpenStudy (mertsj):

50x^2-20x=0 10x(5x-2)=0 10x=0 5x-2=0 x=0, 2/5

OpenStudy (anonymous):

thanks, are you sure? :) I got a different answer

OpenStudy (mertsj):

Well, to begin with, I suspect that you posted the problem incorrectly.

OpenStudy (precal):

use the equation key on the lower left hand side to post your question

OpenStudy (anonymous):

ok jus a sec

OpenStudy (anonymous):

\[50x^3-15x^2-5x=0\]

OpenStudy (precal):

5x(10x^2-3x-1)=0 sorry talking to my boyfriend

OpenStudy (precal):